Prev Next

SW&IT Meeting Minutes 2021-12-13

Attendees: Brandon, David, Glenn, Matthieu, Paul, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20211129
LR20211129 - Bec, Matthieu 80.0
LR20211129 - Mechtley, Brandon 76.0
LR20211129 - Eychaner, Glenn 64.0
LR20211129 - Edmeades, David 64.0
LR20211129 - Grenz, Paul 36.0
LR20211129 - Hooper, Stephen 0.0
LR20211129 - Zhang, Yang 80.0
LR20211129 - Zhang, Xianyu 80.0

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

Team reports

DE

  • Vacation today
  • Success with iSCSI mount on archive-in, will coordinate with all parties to swap newdata to the new volume while keeping the archive software happy
  • VM clone for MB
  • Veeam backups running automatically again
This week:
  • Catchup
  • Mod Veeam config to use local resources
  • Ketiv meeting
  • Test Vault backup script on new versions
  • Followup with AC on the Ethernet ports, stated due date is tomorrow

MB

  • Last week
    • OC, Management
    • Performance Appraisals
    • shark-nir PAE
    • WFS simulator, memleak check
  • This week
    • ATT PM, SHARK-NIR PAE
    • ao-supervisor pyQt3
    • AO PB mini-crate follow-up
    • VAC Dec17 afternoon until Jan

GE

  • Last week
    • Merged the pull request into the LBTO GitHub.
    • Created a release version (2.14.0) of the new LUCI software, and installed it on luci.luci and in stow on the observer machines.
    • Tested the new version of the RTD, and found another serious threading/concurrency issue.
    • Fixed the error handling in the RTD so all errors go to the same place, rather than being spread across System.out, JOptionPanes, and the MessageServer GUI.
  • This week
    • Finish fixing the new threading/concurrency issue in the RTD GUI.
    • Test the new version of the RTD, and release the test version to SciOps.
    • Examine the other LUCI GUIs to check for GUI modifications from threads other than the Event Dispatch Thread

YZ

  • Previous:
    • Several MODS work:
      • Concluded MODS WCS dev/evaluation with Olga (now works very well).
      • MODS IDL evaluation/discussion on new Fedora workstations.
      • A rare MODS Align program glitch is under investigation.
      • MODS operations on obs5/obs6 should be ready, I enabled MODS operations software access for all partner accounts on obs5/obs6.
      • Some IT discussions at the instrument meeting (IT#7400 [concerning MODS TCS service] and IT#8534 [it is really a TCS question]).
    • LBC alert script glitch investigation continues (now I believe it is an email (google group) problem).
    • TCS new leap seconds file downloaded okay (haven't seen any complaints from TCS).
    • Worked a bit on the new image notification utility.
    • TMS meeting and software work (utility update/request, and new impl for missing channel calculation).
  • Next:
    • Conclude/continue some misc MODS work.
    • Fix LBC alert scripts issue.
    • Continue to work on new image notification utility.
    • TMS work continues.
    • LBC bias level wiki page.
    • New safety trainings due.

BM

  • Last week:
    • Updated lbti-sxwfs and sxadsec to latest ao-supervisor. Only soul-sx is behind now, but is tested. Will check in #ao-sw if it's OK to switch.
    • Set up simulation VM more. Blocked communication to mountain servers.
    • Moved all the old sxadsec logs to the mt-synha. The cron job only backs up one day at a time, so it wasn't catching up.
    • Met with MB/GB/XZ about "mixed mode" changes, simulator.
    • Sunday night DEC w/LBTI. Lots of ICE communication errors (communicating SkipFrame alerts) … network troubles? Has only happened 12/7-12/10, 12/13 this year.
  • This week:
    • Vacation T-F.
    • Test config-gain-ramp in simulator.
    • Start working on a lot of smaller, backlogged issues.

PSG

  • Last week
    • Checked in code, made pull request
    • Yang helped fix an issue with checkin - http vs. ssh.
    • Addressing several comments on the pull request.
  • This week
    • Finish implementing changes per comments on pull request.
    • Update pull request.
    • Create schem file for "Jenkins" for Argos Proxy.

SH

  • Increased RAM and CPU on shell for Christian
  • Installed 4K displays on robs
  • Dreamhost stuff
  • VPN issue this weekend
  • UITS, Steward wired registration. I own Steward?
  • Upgrade to FactoryManager for Dan C.
  • Core switch quote sent to orders.

next:

  • Today and tomorrow, then vacation until the end of the year.
  • Some clean up before the EOY and vacation, but I will be callable.
  • ROFIN battery (I will try tomorrow).
~
Topic revision: r10 - 13 Dec 2021, MatthieuBec
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback